# A Large Recurrent Action Model: xLSTM enables Fast Inference for Robotics Tasks

### v1

@@ 1 Introduction
Reinforcement Learning (RL) has been responsible for impressive success stories such as game-playing [Silver et al.,, 2016; Vinyals et al.,, 2019; Berner et al.,, 2019; Patil et al.,, 2022], plasma control for fusion [Degrave et al.,, 2022], or navigation of stratospheric balloons [Bellemare et al.,, 2020]. While these successes were based on classical RL approaches, in which agents have been trained online with RL objectives, recently there has been a trend towards offline RL settings [Levine et al.,, 2020; Schweighofer et al.,, 2022] and sequence models trained via behavior cloning [Chen et al.,, 2021; Janner et al.,, 2021]. Such approaches, in which agents are trained on large-scale offline datasets with causal sequence modeling objectives, have been driven by the proliferation of Transformer-based architectures and gave rise to what we refer to as Large Action Models (LAMs) to highlight their similarity to large language models (LLMs) [Radford et al.,, 2018]. LAM approaches can also be used in multi-task settings to develop generalist agents such as Gato [Reed et al.,, 2022].
@@
Existing LAMs are primarily based on the Transformer [Vaswani et al.,, 2017] architecture. Because of their powerful predictive performance, robotics has become an emergent application area for large models [Brohan et al., 2023b, ; Brohan et al., 2023a, ; Octo Model Team et al.,, 2024; Gu et al.,, 2023; Wang et al.,, 2023] and a number of large multi-task datasets were collected [Jia et al.,, 2024; Embodiment Collaboration et al.,, 2024; Jiang et al.,, 2023; Mandlekar et al.,, 2023]. This development bears the potential to produce robotics agents that learn to master complex tasks in a wide range of environments and even different embodiments. For example, recently it has been demonstrated, albeit in restricted settings, that sequence models trained on multi-episodic contexts can perform in-context learning (ICL) [Laskin et al.,, 2020; Lee et al.,, 2023]. One potential application of ICL can be to learn new related tasks in robotics without the need for re-training or fine-tuning.
One of the key reasons for the success of Transformer-based models is their ability to scale to large datasets through their efficient parallelization during training. However, despite numerous success stories in RL, language modeling [Brown et al.,, 2020] or computer vision [Dosovitskiy et al.,, 2021; He et al.,, 2022], a persistent drawback of Transformer-based architectures is their high inference cost in terms of both speed and memory [Kim et al.,, 2023]. Consequently, deploying Transformer-based models in resource-constrained scenarios, such as on devices with limited hardware capacity and/or real-time constraints, e.g., robots or smartphones, is difficult because of the required fast inference times [Firoozi et al.,, 2023; Hu et al.,, 2023]. A basic principle of control theory is that the controller sample rate should be at least in the order of magnitude of the sample rate of the sensors [Franklin et al.,, 1998, Ch. 11]. To illustrate this, for typical robots such as drones or industrial robot arms, rates of 100Hz-1000Hz are necessary to keep the system stable [Salzmann et al.,, 2023; El-Hussieny,, 2024; Hu et al.,, 2023; Chignoli et al.,, 2021]. This implies inference times of less than 10ms. At 1000Hz, a 15-second movement of the agent corresponds to a sequence of 15K steps [El-Hussieny,, 2024] resulting in long context lengths even without ICL. While there exists a range of techniques to make large models faster, such as quantization [Frantar et al.,, 2023], distillation [Hinton et al.,, 2015], or pruning [LeCun et al.,, 1989], the quadratic-time complexity of self attention still remains.
Recently, modern recurrent architectures have been proposed, which exhibit similar parallelization properties during training as the Transformer architecture while offering linear-time inference complexity. These modern recurrent architectures include xLSTM [Beck et al.,, 2024] and state-space models (SSMs), such as Mamba [Gu and Dao,, 2023; Dao and Gu,, 2024] and Griffin/Hawk [De et al.,, 2024], and have challenged the dominance of the Transformer in language modeling but also in other domains such as computer vision [Alkin et al.,, 2024; Zhu et al.,, 2024], and biomedicine [Schmidinger et al.,, 2024]. More importantly, their linear-time inference makes them suitable for deployment in scenarios with limited compute, large context sizes, and real-time requirements, such as robotics.
@@
Figure 1: Illustration of our Large Recurrent Action Model (LRAM) with an xLSTM [Beck et al.,, 2024] at its core.
@@
In this work, we assess the aptitude of modern recurrent architectures, such as xLSTM and Mamba, as large action models. To this end, we introduce a Large Recurrent Action Model (LRAM) with an xLSTM at its core (see Figure 1). We train our agents on 432 tasks from 6 domains using a supervised learning setting similar to that of the Decision Transformer [Chen et al.,, 2021, DT]. We use data collected during online-RL training of single-task specialist agents and compile these trajectories alongside other expert demonstrations into a large-scale multi-domain dataset comprising 894M transitions. Due to their parallelization properties, the modern recurrent architectures considered in this work can process this large-scale training set as efficiently as the Transformer while being faster at inference. Experiments across 4 models sizes with our multi-task models indicate that xLSTM compares favorably to Transformers in terms of both performance and speed. In addition, we study the effect of modern recurrent architectures on fine-tuning performance and in-context learning abilities, and find that they exhibit strong performance in both dimensions.
@@
  • To foster further research on large action models, we release our data preparation pipeline and generated datasets [^1] GitHub: https://github.com/ml-jku/LRAM.
@@ 2 Related work
Sequence Models in RL. LSTM [Hochreiter and Schmidhuber,, 1997] is the dominant backbone architecture for partially observable online RL problems and has been behind achievements such as mastering Starcraft II [Vinyals et al.,, 2019], Dota 2 [Berner et al.,, 2019], and Atari [Espeholt et al.,, 2018; Kapturowski et al.,, 2019]. After the success of the Transformer in NLP [Devlin et al.,, 2019; Radford et al.,, 2019; Brown et al.,, 2020], computer vision [Dosovitskiy et al.,, 2021; He et al.,, 2022; Radford et al.,, 2021; Fürst et al.,, 2022] and speech recognition [Radford et al.,, 2022; Baevski et al.,, 2020], the architecture has found its way into RL. Chen et al., [2021] proposed the Decision Transformer (DT) a GPT-style model [Radford et al.,, 2018], that learns to predict actions from offline trajectories via behavior cloning. Trajectory Transformer [Janner et al.,, 2021] predicts action tokens along with states and rewards, which allows for dynamics modeling. A number of follow-up works build on the DT-architecture [Zheng et al.,, 2022; Wang et al.,, 2022; Shang et al.,, 2022; Meng et al.,, 2021; Siebenborn et al.,, 2022; Schmied et al., 2024a, ]. Furthermore, sequence models trained on offline data were found to exhibit in-context learning if conditioned on previous trajectories [Laskin et al.,, 2022; Lee et al.,, 2022; Kirsch et al.,, 2023], albeit only in limited scenarios.
@@
Large Action Models (LAMs). LAMs, such as the Decision Transformer, are well suited for multi-task settings. Lee et al., [2022] found that a multi-game DT can learn to play 46 Atari games. Reed et al., [2022] introduced a generalist agent trained on over 600 tasks from different domains, ranging from Atari to manipulation of a robot arm. Jiang et al., [2022] a Transformer for robot manipulation based on multi-modal prompts, that allow to steer the model to perform new tasks. Recently, Raad et al., [2024] introduced an agent instructable via language to play a variety of commercial video games. Since then, robotics has become an emergent area for developing LAMs [Brohan et al., 2023b, ; Brohan et al., 2023a, ; Octo Model Team et al.,, 2024; Gu et al.,, 2023; Wang et al.,, 2023; Di Palo et al.,, 2023; Kim et al.,, 2024], also due to the availability of large-scale robotics datasets [Jia et al.,, 2024; Embodiment Collaboration et al.,, 2024; Jiang et al.,, 2023; Mandlekar et al.,, 2023].
@@
Next-generation Sequence Modeling Architectures. Linear recurrent models, such as state-space models (SSM, Gu et al.,, 2021; Gu et al., 2022b, ; Smith et al.,, 2023; Orvieto et al.,, 2023) have challenged the dominance of the Transformer [Vaswani et al.,, 2017] architecture on long-range tasks [Tay et al.,, 2020]. The key insight of those linear RNNs was to diagonalize the recurrent state matrix and enforce stable training via an exponential parameterization [Gu et al., 2022a, ; Orvieto et al.,, 2023]. Since then, there have been efforts to include features such as gating from RNNs [Elman,, 1990; Jordan,, 1990; Hochreiter and Schmidhuber,, 1997; Cho et al.,, 2014]. Non-linear gates are believed to have higher expressivity, but are harder to train. Griffin [De et al.,, 2024] mixes gated linear recurrences with local attention to achieve more training data efficiency than Llama-2 [Touvron et al.,, 2023] and better sequence extrapolation. Mamba [Gu and Dao,, 2023] introduces a selection mechanism similar to gating into SSMs, which makes its state and input matrix time dependent. This is similar to the gating mechanism of RNNs but also bears resemblance to approaches like fast weights [Schmidhuber,, 1992] and Linear Attention [Katharopoulos et al.,, 2020]. Mamba-2 [Dao and Gu,, 2024] highlight the connection between SSMs with input dependent state and input matrices and (Gated) Linear attention variants. Most recently, the xLSTM [Beck et al.,, 2024] was proposed as an improvement over the classic LSTM [Hochreiter and Schmidhuber,, 1997] that combines gating, linear recurrences and recurrent weights into a single architecture for language modeling. First, xLSTM leverages exponential gating with stabilization to RNNs for stronger emphasis on important inputs. Second, xLSTM is composed of two variants, the mLSTM variant with an emphasis on memory that proves important in language modeling and the sLSTM variant that keeps the non-diagonalized recurrent matrix to enable state-tracking [Merrill et al.,, 2024]. State tracking is important in logic tasks and cannot be modeled fundamentally by linearized recurrent or state-space models like Mamba, Griffin or Transformers.
3 Large Recurrent Action Models
3.1 Background
Decision Transformer [Chen et al.,, 2021] casts the RL problem setting as next action prediction task via causal sequence modeling. At training time, DT aims to learn a policy $\pi_{\theta}$ that maps future rewards to actions, which is often referred to as upside-down RL [Schmidhuber,, 2019]. At inference time, the DT is conditioned via a target return to emit high-reward actions. Consequently, we assume access to a dataset $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$ containing $N$ trajectories $\tau_{i}$ consisting of quadruplets $\tau_{i}=(s_{1},\hat{R}_{1},a_{1},r_{1},\dots,s_{T},\hat{R}_{T},a_{T},r_{T})$ of state $s_{t}$, return-to-go (RTG) $\hat{R}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$, action $a_{t}$, and reward $r_{t}$. Here, $T$ refers to the length of the trajectory. The DT $\pi_{\theta}$ is trained to predict the ground-truth action $a_{t}$ conditioned on sub-trajectories from the dataset:
$\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}\mid s_{t-C},\hat{R}_{t-C},a_{t-C},r_{t-C},\dots,s_{t-1},\hat{R}_{t-1},a_{t-1},r_{t-1},s_{t},\hat{R}_{t}),$ (1)
where $C\leq T$ is the size of the context window. In fact, Equation 1 describes the setting of the multi-game DT [Lee et al.,, 2022], which also includes rewards in the sequence representation.
3.2 Large Recurrent Action Models (LRAMs)
Our LRAM has a modern recurrent architecture at its core (see Figure 1), which comes with a parallel training and a recurrent inference mode. We instantiate LRAM with three different variants, two different xLSTM configurations and Mamba. Furthermore, we use a training protocol similar to that of Lee et al., [2022] and Reed et al., [2022] with some differences.
Multi-modal sequence representation. To encode input from different environments with varying state and action spaces, we use separate encoders per modality that are shared across tasks and domains. For encoding images we use a CNN similar to Espeholt et al., [2018], whereas for low-dimensional inputs we use a fully connected network. We refrain from patchifying images and tokenizing continuous states to avoid unnecessarily long sequences. Similarly, we use linear layers to encode rewards and RTGs. We omit actions in our sequence formulation, as we found that this can be detrimental to performance, in particular for continuous control tasks (see Section 4.3). Consequently, our trajectories have the form $\tau_{i}=(s_{1},\hat{R}_{1},r_{1},\dots,s_{T},\hat{R}_{T},r_{T})$ and we train our policy $\pi_{\rho}$ to predict the ground-truth action $a_{t}$ as:
Shared action head. Action spaces in RL typically vary across environments. For example, in the environments we consider, there are 18 discrete actions and a maximum of 8 continuous dimensions for continuous control environments. Therefore, we employ discretization of continuous action dimensions into 256 uniformly-spaced bins, similar to Reed et al., [2022] and Brohan et al., 2023b . Unlike prior work, we leverage a shared action head to predict all discrete actions or continuous action dimensions jointly. We found this setup significantly reduces inference time compared to using autoregressive action prediction of continuous actions.
Recurrent inference mode. At inference time, we leverage the recurrent backbone and maintain the hidden states of the last timestep. This enables fast inference with linear-time complexity along the sequence length. In addition, the recurrent-style inference is well suited for online fine-tuning via RL objectives, similar to LSTM-based policies in online RL. To further speed-up inference, we leverage custom kernels for the xLSTM backbone (see Appendix 22).
Our unified discrete action representation enables consistent training of our agents via the cross-entropy loss as training objective across all tasks and domains, similar to Reed et al., [2022]. We use separate reward scales per domain and target returns per task. Furthermore, we do not make use of timestep encodings as used by Chen et al., [2021], which are detrimental when episode lengths vary. We provide additional implementation details in Appendix B.
4 Experiments
We study the aptitude of modern recurrent architectures as LAMs on 432 tasks from 6 domains: Atari [Bellemare et al.,, 2013], Composuite [Mendez et al.,, 2022], DMControl [Tassa et al.,, 2018], Meta-World [Yu et al., 2020b, ], Mimicgen [Mandlekar et al.,, 2023], and Procgen [Cobbe et al., 2020a, ]. To this end, we compile a large-scale dataset containing 894 million transitions (see Section 4.1).
Across all experiments, we compare four backbone variants: xLSTM [7:1], xLSTM [1:0] [Beck et al.,, 2024], Mamba [Gu and Dao,, 2023], and the GPT-2 style Transformer employed in the DT [Chen et al.,, 2021]. Following [Beck et al.,, 2024], we use the bracket notation for xLSTM, which indicates the ratio of mLSTM to sLSTM blocks. For example, xLSTM [1:0] contains only mLSTM blocks.
In Section 4.2, we conduct a scaling comparison for four model sizes ranging from 16M to 208M parameters that shows that modern recurrent architectures achieve performance comparable or favorable to the Transformer baseline across different model sizes. In Section 4.3, we study the impact of the recurrent backbones on fine-tuning performance and ICL abilities, and further analyze our trained recurrent backbones. Finally, in Section 4.4, we empirically examine the differences at inference time in terms of latency and throughput between xLSTM-based and Transformer-based agents, which indicate a clear advantage for the recurrent backbone.
4.1 Datasets & Environments
@@
Datasets. We compile a large-scale dataset comprising 432 tasks from six domains. We leverage datasets from prior works. For Atari, we extract 5M transitions per task from the DQN-Replay dataset released by Agarwal et al., [2020]. For Composuite, we leverage the datasets released by [Hussing et al.,, 2023]. For Meta-World, we use 2M transitions per task released by [Schmied et al., 2024a, ]. For DMControl, we generate 10M transitions per task using task-specific RL agents. For Mimicgen, we use the datasets for the 21 tasks released by [Mandlekar et al.,, 2023] and generate trajectories for the remaining 62 tasks. Finally, for Procgen, we extract 20M transitions from the datasets released by [Schmied et al., 2024b, ]. Our final dataset contains 3.4M trajectories and in total 894M transitions (see Table 1). We reserve an additional 37 tasks from the same domains for zero-shot evaluation. To foster future research, we release our data-preparation pipeline and generated data.
Environments. Atari and Procgen come with image observations and discrete actions. In contrast, the remaining four domains exhibit state-based observations and continuous actions. Consequently, our experiments involve a mixture of state and action spaces as well as varying episode lengths (see Table 1). Periodically evaluating the trained agents on all 432 tasks sequentially is time-consuming and we, therefore, distributed the evaluation across GPUs and parallel processes (see Appendix B).
Additional details on our datasets, environments are available in Appendix A.

### v2

@@ 1 Introduction
Reinforcement Learning (RL) has been responsible for impressive success stories such as game-playing (Silver et al., 2016; Vinyals et al., 2019; Berner et al., 2019; Patil et al., 2022), plasma control for fusion (Degrave et al., 2022), or navigation of stratospheric balloons (Bellemare et al., 2020). While these successes were based on classical RL approaches, in which agents have been trained online with RL objectives, recently there has been a trend towards offline RL settings (Levine et al., 2020; Schweighofer et al., 2022) and sequence models trained via behavior cloning (Chen et al., 2021; Janner et al., 2021). Such approaches, in which agents are trained on large-scale offline datasets with causal sequence modeling objectives, have been driven by the proliferation of Transformer-based architectures and gave rise to what we refer to as Large Action Models (LAMs) to highlight their similarity to large language models (LLMs) (Radford et al., 2018). LAM approaches can also be used in multi-task settings to develop generalist agents such as Gato (Reed et al., 2022).
@@
Existing LAMs are primarily based on the Transformer (Vaswani et al., 2017) architecture. Because of their powerful predictive performance, robotics has become an emergent application area for large models (Brohan et al., 2023b, a; Octo Model Team et al., 2024; Gu et al., 2023; Wang et al., 2023) and a number of large multi-task datasets were collected (Jia et al., 2024; Embodiment Collaboration et al., 2024; Jiang et al., 2023; Mandlekar et al., 2023). This development bears the potential to produce robotics agents that learn to master complex tasks in a wide range of environments and even different embodiments. For example, recently it has been demonstrated, albeit in restricted settings, that sequence models trained on multi-episodic contexts can perform in-context learning (ICL) (Laskin et al., 2020; Lee et al., 2023). One potential application of ICL can be to learn new related tasks in robotics without the need for re-training or fine-tuning.
@@
Figure 1: Illustration of our Large Recurrent Action Model (LRAM) with an xLSTM (Beck et al., 2024) at its core.
One of the key reasons for the success of Transformer-based models is their ability to scale to large datasets through their efficient parallelization during training. However, despite numerous success stories in RL, language modeling (Brown et al., 2020) or computer vision (Dosovitskiy et al., 2021; He et al., 2022), a persistent drawback of Transformer-based architectures is their high inference cost in terms of both speed and memory (Kim et al., 2023). Consequently, deploying Transformer-based models in resource-constrained scenarios, such as on devices with limited hardware capacity and/or real-time constraints, e.g., robots or smartphones, is prohibitive because of the required fast inference times (Firoozi et al., 2023; Hu et al., 2023). A basic principle of control theory is that the controller sample rate should be in the order of magnitude of the sample rate of the sensors (Franklin et al., 1998, Ch. 11). To illustrate this, for typical robots such as drones or industrial robot arms rates of 100Hz-1000Hz are required to keep the system stable (Salzmann et al., 2023; El-Hussieny, 2024; Hu et al., 2023; Chignoli et al., 2021). This implies inference times of less than 10ms. At 1000Hz, a 15-second movement of the agent corresponds to a sequence of 15K steps (El-Hussieny, 2024) resulting in long context lengths even without ICL. While there exists a range of techniques to make large models faster, such as quantization (Frantar et al., 2023), distillation (Hinton et al., 2015), or pruning (LeCun et al., 1989), the quadratic-time complexity of self attention still remains.
Recently, modern recurrent architectures have been proposed, which exhibit similar parallelization properties during training as the Transformer architecture while offering linear-time inference complexity. These modern recurrent architectures include xLSTM (Beck et al., 2024) and state-space models (SSMs), such as Mamba (Gu & Dao, 2023; Dao & Gu, 2024) and Griffin/Hawk (De et al., 2024), and have challenged the dominance of the Transformer in language modeling but also in other domains such as computer vision (Alkin et al., 2024; Zhu et al., 2024), and biomedicine (Schmidinger et al., 2024). More importantly, their linear-time inference makes them suitable for deployment in scenarios with limited compute, large context sizes, and real-time requirements, such as robotics.
@@
In this work, we assess the aptitude of modern recurrent architectures, such as xLSTM and Mamba, as large action models. To this end, we introduce a Large Recurrent Action Model (LRAM) with an xLSTM at its core (see Figure 1). We train our agents on 432 tasks from 6 domains using a supervised learning setting similar to that of the Decision Transformer (Chen et al., 2021, DT). We use data collected during online-RL training of single-task specialist agents and compile these trajectories alongside other expert demonstrations into a large-scale multi-domain dataset comprising 894M transitions. Due to their parallelization properties, the modern recurrent architectures considered in this work can process this large-scale training set as efficiently as the Transformer while being faster at inference. Experiments across 4 models sizes with our multi-task models indicate that xLSTM compares favorably to Transformers in terms of both performance and speed. In addition, we study the effect of modern recurrent architectures on fine-tuning performance and in-context learning abilities, and find that they exhibit strong performance in both dimensions.
@@
  • To foster further research on large action models, we release our data preparation pipeline and our datasets. [^1] GitHub: https://github.com/ml-jku/LRAM
@@ 2 Related work
Sequence Models in RL. LSTM (Hochreiter & Schmidhuber, 1997) is the dominant backbone architecture for partially observable online RL problems and has been behind achievements such as mastering Starcraft II (Vinyals et al., 2019), Dota 2 (Berner et al., 2019), and Atari (Espeholt et al., 2018; Kapturowski et al., 2019). After the success of the Transformer in NLP (Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; He et al., 2022; Radford et al., 2021; Fürst et al., 2022) and speech recognition (Radford et al., 2022; Baevski et al., 2020), the architecture has found its way into RL. Chen et al. (2021) proposed the Decision Transformer (DT) a GPT-style model (Radford et al., 2018), that learns to predict actions from offline trajectories via behavior cloning. Trajectory Transformer (Janner et al., 2021) predicts actions along with states and rewards, which allows for dynamics modeling. Other follow-up works build on the DTs (Zheng et al., 2022; Wang et al., 2022; Shang et al., 2022; Meng et al., 2021; Siebenborn et al., 2022; Schmied et al., 2024a) or replace the Transformer with Mamba (Ota, 2024; Dai et al., 2024). Furthermore, sequence models trained were found to exhibit ICL if conditioned on previous trajectories (Laskin et al., 2022; Lee et al., 2022; Kirsch et al., 2023), albeit in limited scenarios.
@@
Large Action Models (LAMs). LAMs, such as the Decision Transformer, are well suited for multi-task settings. Lee et al. (2022) found that a multi-game DT can learn to play 46 Atari games. Reed et al. (2022) introduced a generalist agent trained on over 600 tasks from different domains, ranging from Atari to manipulation of a robot arm. Jiang et al. (2022) a Transformer for robot manipulation based on multi-modal prompts, that allow to steer the model to perform new tasks. Recently, Raad et al. (2024) introduced an agent instructable via language to play a variety of commercial video games. Since then, robotics has become an emergent area for developing LAMs (Brohan et al., 2023b, a; Octo Model Team et al., 2024; Gu et al., 2023; Wang et al., 2023; Kim et al., 2024), also due to the availability of large-scale datasets (Jia et al., 2024; Embodiment Collaboration et al., 2024; Jiang et al., 2023; Mandlekar et al., 2023).
@@
Next-generation Sequence Modeling Architectures. Linear recurrent models, such as state-space models (SSM, Gu et al., 2021, 2022b; Smith et al., 2023; Orvieto et al., 2023) have challenged the dominance of the Transformer (Vaswani et al., 2017) architecture on long-range tasks (Tay et al., 2020). The key insight of those linear RNNs was to diagonalize the recurrent state matrix and enforce stable training via an exponential parameterization (Gu et al., 2022a; Orvieto et al., 2023). Since then, there have been efforts to include features such as gating from RNNs (Elman, 1990; Jordan, 1990; Hochreiter & Schmidhuber, 1997; Cho et al., 2014). Non-linear gates are believed to have higher expressivity, but are harder to train. Griffin (De et al., 2024) mixes gated linear recurrences with local attention to achieve more training data efficiency than Llama-2 (Touvron et al., 2023) and better sequence extrapolation. Mamba (Gu & Dao, 2023) introduces a selection mechanism similar to gating into SSMs, which makes its state and input matrix time dependent. This is similar to the gating mechanism of RNNs but also bears resemblance to approaches like fast weights (Schmidhuber, 1992) and Linear Attention (Katharopoulos et al., 2020). Mamba-2 (Dao & Gu, 2024) highlight the connection between SSMs with input dependent state and input matrices and (Gated) Linear attention variants. Most recently, the xLSTM (Beck et al., 2024) was proposed as an improvement over the classic LSTM (Hochreiter & Schmidhuber, 1997) that combines gating, linear recurrences and recurrent weights into a single architecture for language modeling. First, xLSTM leverages exponential gating with stabilization to RNNs for stronger emphasis on important inputs. Second, xLSTM is composed of two variants, the mLSTM variant with an emphasis on memory that proves important in language modeling and the sLSTM variant that keeps the non-diagonalized recurrent matrix to enable state-tracking (Merrill et al., 2024). State tracking is important in logic tasks and cannot be modeled fundamentally by linearized recurrent or state-space models like Mamba, Griffin or Transformers.
@@
3 Large Recurrent Action Models
3.1 Background
Decision Transformer (Chen et al., 2021) casts the RL problem setting as next action prediction task via causal sequence modeling. At training time, DT aims to learn a policy $\pi_{\theta}$ that maps future rewards to actions, which is often referred to as upside-down RL (Schmidhuber, 2019). At inference time, the DT is conditioned via a target return to emit high-reward actions. Consequently, we assume access to a dataset $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$ containing $N$ trajectories $\tau_{i}$ consisting of quadruplets $\tau_{i}=(s_{1},\hat{R}_{1},a_{1},r_{1},\dots,s_{T},\hat{R}_{T},a_{T},r_{T})$ of state $s_{t}$, return-to-go (RTG) $\hat{R}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$, action $a_{t}$, and reward $r_{t}$. Here, $T$ refers to the length of the trajectory. The DT $\pi_{\theta}$ is trained to predict the ground-truth action $a_{t}$ conditioned on sub-trajectories from the dataset:
$\begin{split}\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}\mid&s_{t-C},\hat{R}_{t-C},a_{t-C},r_{t-C},\dots,\\ &s_{t-1},\hat{R}_{t-1},a_{t-1},r_{t-1},s_{t},\hat{R}_{t})\end{split}$ (1)
where $C\leq T$ is the size of the context window. In fact, Equation 1 describes the setting of the multi-game DT (Lee et al., 2022), which also includes rewards in the sequence representation.
3.2 Large Recurrent Action Models (LRAMs)
Our LRAM has a modern recurrent architecture at its core (see Figure 1), which comes with a parallel training and a recurrent inference mode. We instantiate LRAM with three different variants, two different xLSTM configurations and Mamba. We use a training protocol similar to that of Lee et al. (2022) and Reed et al. (2022) with important differences that aim to speed-up inference across backbones.
Multi-modal sequence representation. To encode input from different environments with varying state and action spaces, we use separate encoders per modality that are shared across tasks and domains. For encoding images we use a CNN similar to Espeholt et al. (2018), whereas for low-dimensional inputs we use a fully connected network. We refrain from patchifying images and tokenizing continuous states to avoid unnecessarily long sequences. Similarly, we use linear layers to encode rewards and RTGs. We omit actions in our sequence formulation, as we found that this can be detrimental to performance, in particular for continuous control tasks with smoothly changing actions (see Section 4.3). Consequently, our trajectories have the form $\tau_{i}=(s_{1},\hat{R}_{1},r_{1},\dots,s_{T},\hat{R}_{T},r_{T})$ and we train our policy $\pi_{\rho}$ to predict the ground-truth action $a_{t}$ as:
Shared action head. Action spaces in RL typically vary across environments. For example, in the environments we consider, there are 18 discrete actions and a maximum of 8 continuous dimensions for continuous control environments. Therefore, we employ discretization of continuous action dimensions into 256 uniformly-spaced bins, similar to Reed et al. (2022) and Brohan et al. (2023b). Unlike prior work, we leverage a shared action head to predict all discrete actions or continuous action dimensions at jointly. We found this setup significantly reduces inference time compared to using autoregressive action prediction of continuous actions.
Recurrent inference mode. At inference time, we leverage the recurrent backbone and maintain the hidden states of the last timestep. This enables fast inference with linear-time complexity along the sequence length. In addition, the recurrent-style inference is well suited for online fine-tuning via RL objectives, similar to LSTM-based policies in online RL. To speed-up inference, we leverage custom kernels for the xLSTM backbone (see Appendix 21).
Our unified discrete action representation enables consistent training of our agents via the cross-entropy loss as training objective across all tasks and domains, similar to Reed et al. (2022). We use separate reward scales per domain and target returns per task. Furthermore, we do not make use of timestep encodings as used by Chen et al. (2021), which are detrimental when episode lengths vary. We provide additional implementation details in Appendix C.